# The Interaction Layer: An Exploration for Co-Designing User-LLM Interactions in Parental Wellbeing Support Systems

### v1

@@ Abstract.
Parenting brings emotional and physical challenges, from balancing work, childcare, and finances to coping with exhaustion and limited personal time. Yet, one in three parents never seek support. AI systems potentially offer stigma-free, accessible, and affordable solutions. Yet, user adoption often fails due to issues with explainability and reliability. To see if these issues could be solved using a co-design approach, we developed and tested NurtureBot, a wellbeing support assistant for new parents. 32 parents co-designed the system through Asynchronous Remote Communities method, identifying the key challenge as achieving a “successful chat.” As part of co-design, parents role-played as NurturBot, rewriting its dialogues to improve user understanding, control, and outcomes. The refined prototype evaluated by 32 initial and 46 new parents, showed improved user experience and usability, with final CUQ score of 91.3/100, demonstrating successful interaction patterns. Our process revealed useful interaction design lessons for effective AI parenting support.
Figure 1. Overview of the six-part study detailing the iterative development and testing of NurtureBot, a parental wellbeing assistant.
@@
A visual representation of the six-part iterative development of NurtureBot involving two parent cohorts. The chart shows the progression from the initial prototype to NurtureBot v1, followed by ARC co-design sessions, NurtureBot v2, and NurtureBot v3. Key findings include conversation starters, tone, parenting resources, and interaction improvements. NurtureBot v1 identified seven major problems, leading to an interaction layer for NurtureBot v2, which addressed four feature requests. NurtureBot v3 showed further improvements with five new feature requests and achieved the highest CUQ score of 91.3.
@@ 1. Introduction
Parenting, often hailed as the most important job in our world, exacts a significant toll on both mind and body. The perinatal period represents a critical developmental stage for the child, the parent(s), and the formation of attachment relationships (Li et al., 2017; Allan et al., 2013). This period, spanning from the onset of pregnancy to the first year postpartum, is marked by profound physical, emotional, and psychological changes, which significantly impact parental wellbeing. In particular, many expectant and new mothers encounter substantial challenges, in mental health, emotional adjustment, and caregiving (Widarsson et al., 2012), in addition to the transformation their body goes through. During this period, parents experience heightened stress, anxiety, and feelings of isolation, increasing the need for support in these areas. Yet, we know that one in three parents do not seek help (UNICEF UK, 2022). The literature on perinatal care identifies several key issues, including inadequate access to mental health resources, availability of healthcare professionals, cost to the government, stigma surrounding parental mental health, and challenges in identifying and supporting at-risk mothers (Rallis et al., 2014; Davis et al., 2022; Howard and Khalifeh, 2020). These challenges have been further aggravated by the COVID-19 pandemic (Bridle et al., 2022; Kinser et al., 2022). As a result, we see a growing body of research exploring how digital technologies can enhance perinatal care and provide support during this critical period (van den Heuvel et al., 2018; Ginja et al., 2018).
@@
Notable Human-Computer Interaction (HCI) research has laid the groundwork for exploring how digital interventions increase wellbeing across aspects such as physical and mental health of a family (Pina et al., 2017), a child’s growth (Kientz et al., 2009), sleep tracking of parents and children (Pina et al., 2020), and managing family screen-time behaviours (Hiniker et al., 2016). Today, with the availability of Artificial Intelligence (AI) to everyday users, use of general generative applications such a ChatGPT for parenting is on the rise (Quan et al., 2024). With responsible design, Large Language Models (LLMs) have the potential to facilitate a range of services, from offering self-guided mental health support (Sharma et al., 2024), real-time emotional support (Kang et al., 2024) to providing evidence-based information at scale (Lai et al., 2023). The current use of LLMs to support parental and child wellbeing is reviewed with high potential and ethical cautions (Ashraf, 2024; Stade et al., 2024). Their application within this context is fraught with limitations and uncertainties, especially of accuracy, empathy, and usefulness (Kaneda et al., 2023).
@@
Given these challenges, our research adopts an exploratory approach, probing into a small subset of these broader issues. We focus on understanding how interactions between LLMs and parents might be designed to be perceived as helpful and engaging. Our goal is to explore what constitutes a usable and useful interaction in these contexts and how LLMs can be integrated into wellbeing support in ways that are practically engaging and thus useful for parents. To this end, we conducted a six-part online study involving 78 parents, who trialled iterative improvements of NurtureBot, a technology probe, designed to support parental wellbeing. This study, while not aimed at delivering a finalised solution, served as an exploratory step toward understanding how parents interact with LLM-based agents and identifying necessary features to make these tools genuinely supportive. The following Research Questions (RQs) guide our investigation into the user interaction aspects of engaging with an LLM-based parental wellbeing assistant:
@@
This six-part user-centred design process, as illustrated in Figure 1, guided the iterative development of NurtureBot. To mitigate emotional risks, we engaged a ‘slightly-out-of-target’ group of parents, each with a child aged five or six. These parents simulated stages of their perinatal period and reflected on past challenges with NurtureBot. Using a mixed-methods approach, we collected both qualitative and quantitative feedback, with NurtureBot v1, v2, and v3 tested over a five-day period.On the final day, participants completed the Chatbot Usability Questionnaire (CUQ) (Holmes et al., 2019), assessing usability and engagement. (see Figure2).
@@
  • Part I: The Initial Prototype: NurtureBot v1 The development of NurtureBot v1 leveraged LLM technology, informed by a community wellbeing initiative supporting new parents. Zero-shot prompting was used to integrate supportive conversation starters and key features, including empathetic chatting, wellbeing exercises, and parenting information (see Section 3).
  • Part II: NurtureBot v1 Testing and Analysis A cohort of 32 parents engaged with NurtureBot v1, generating 144 conversations. Thematic analysis identified several issues, including lack of conversational flow, robotic tone, transactional attitude, limited personalisation, and localisation. Despite these challenges, NurtureBot v1 achieved a CUQ score of 85.4, surpassing the benchmark of 68 (Holmes et al., 2019), indicating good usability (see Section 4).
@@
  • Part III: ARC Co-Design Sessions and Analysis The Asynchronous Remote Community (ARC) (MacLeod et al., 2016; Prabhakar et al., 2017) co-design sessions engaged the same cohort of parents with Miro board activities (Miro, 2011).
    • Parents first prioritised key issues, with “successful chat” emerging as the top priority.
    • Then they reimagined and rewrote NurtureBot’s dialogues, aiming for better user understanding, user control, and improved outcomes, a structure akin to a cybernetic loop (Patten and Odum, 1981).
    • Thematic analysis of 189 revised dialogues led to requests for essential features such as metaphorical explanations, empathetic engagement, feature introductions, personalising exchanges, and step-by-step guidance (see Section 5).
@@
  • Part IV: The Interaction Layer: NurtureBot v2 NurtureBot v2 introduced an interaction layer with three essential components (See Figure 7): (i) A core interaction principle that enabled users to ”Understand, Control, and Improve” exchanges, (ii) Interaction levels based on few-shot prompting, incorporating co-designed dialogues from ARC sessions, and (iii) Interaction elements, tailored to address user needs identified during ARC.
  • Part V: NurtureBot v2 Testing and Analysis NurtureBot v2 underwent testing with the same initial cohort, yielding 104 conversations. Our analysis revealed that two minor issues persisted: repetitive conversational hints and the delivery of parenting information in long, overwhelming paragraphs. Additionally, themes such as localisation and conversational memory, which we did not prioritise addressing in this study, continued to emerge. NurtureBot v2 achieved a CUQ score of 87.4, indicating further improvement, though the potential bias from participants’ involvement in co-design was considered (see Section 7).
@@
  • Part VI: Prototype Improvement: NurtureBot v3 and Validation To ensure an unbiased evaluation, we refined NurtureBot further and tested v3 with a new cohort of 46 parents. Feedback analysis identified two new problems, the need for more detailed responses at times, difficulty following step-by-step exercises via text, and repeated requests for additional features such as videos and memory retention for personalised interactions. NurtureBot v3 received a CUQ score of 91.3, marking a robust improvement in usability (see Section 8).
@@
This study initiates our exploration into co-designing agents for parental wellbeing, setting the foundation for future research to refine and expand AI tools. The diverse parenting experiences and emotional complexities highlight the need for continued development and testing.
@@ 2.1. Supporting Family Wellbeing with Digital Tools
Within the field of HCI, there has been a long-term research agenda focusing on the study and development of impactful digital tools that offer a variety of social, physical, and mental wellbeing support to mothers (D’Ignazio et al., 2016; Thomas et al., 2019; Britton et al., 2019). The design and evaluation of digital tools aimed at promoting maternal wellbeing during the transition to motherhood (Newhouse, 2019), bump2bump, emphasise the need for holistic approaches in HCI that consider the challenges faced by new mothers. The role of peer support in digital environments is highlighted by the finding that online networks can provide valuable emotional and practical assistance to parents, particularly in situations where traditional face-to-face support may not be available (Yamashita et al., 2022). The use of the Baby Buddy app among first-time mothers in the UK (Deave et al., 2019), shows that while the app is generally well-received, its effectiveness is influenced by how well it aligns with the mothers’ understanding of their needs and the app’s usability. This points to the necessity of involving parents in the design process to ensure that digital tools meet their expectations and provide meaningful support.
@@
The integration of AI chatbots into parenting and family well-being has opened new avenues for providing support and guidance to parents (Yu et al., 2023; Entenberg et al., 2021; Wong et al., 2021). Parenting with conversational agents such as Alexa (Beneteau et al., 2020), Siri (Sezgin et al., 2022) and other voice-activated agents in the home (Sun et al., 2021) have been critically studied in the past few years. Following the advent of AI chatbots, the integration of LLMs into parenting practices (Quan et al., 2024), is ushering in an era of unprecedented possibilities and challenges. LLM-based agents such as ChatGPT are being adopted by parents of young children (Quan et al., 2024; Kaneda et al., 2023), offering new learning opportunities and personalised support, though their impact on parenting remains largely unexplored, marking this as an era of the unknown. The performance and acceptance of ChatGPT’s responses in the childcare field (Kaneda et al., 2023), is also being studied with caution. Despite technological advancements, several limitations persists such as misinformation (Kaneda et al., 2023), lack of diversity (Bajwa et al., 2024), threat of data misuse (Neerincx et al., 2021), and lack of user adoption (Baumel et al., 2019). While gamification can be an effective motivational tool (Wernbacher et al., 2022), it often does not appeal to all users equally, particularly those unfamiliar with or resistant to digital game mechanics. The need for humane interaction with a good user experience has been repeatedly reported in the literature for digital health tools built with or without AI (Bautista and Schueller, 2023). A review of 13 AI-based mental health apps with user ratings over 4 out of 5, revealed limited support to AI literacy and explainability (Alotaibi and Sas, 2023). Real-world engagement data from 93 health apps indicate that while install rates are high, sustained user engagement remains low, with the median 15- and 30-day retention a mere 3.9% and 3.3% respectively (Baumel et al., 2019). These gaps suggest a need for further research and development of ‘truly interactive’ applications to deliver sound AI-powered digital tools for wellbeing.
@@ 2.2. Interacting with AI for Everyday Users
As AI permeates daily life, from recommendation systems to digital assistants, the ability of non-expert users to engage meaningfully with AI systems becomes critical (Eiband et al., 2019). Research highlights the significance of Explainable AI (XAI) in demystifying the opaque processes of machine learning models, thereby enhancing user understanding, an extensively studied dimension of Human-Centred AI (HCAI)(Saeed and Omlin, 2023; Capel and Brereton, 2023). XAI methods such as Local Interpretable Model-agnostic Explanations (LIME) and SHapley Additive exPlanations (SHAP) have been pivotal in rendering AI decision-making processes more transparent (Ribeiro et al., 2016; Lundberg and Lee, 2017; Adadi and Berrada, 2018; Arrieta et al., 2020). While XAI tools enable expert users to grasp the rationale behind AI outputs, (Gunning et al., 2019; Miller, 2023a; Carvalho et al., 2019), most non-experts do not engage with XAI (Miller, 2023b), highlighting the need for innovative methods to foster AI understanding (Miller, 2023b). Control over AI systems by everyday users is another essential dimension of HCAI, with research underscoring the importance of user agency in human-AI interactions and challenges with it (König, 2024). Interactive interfaces that allow users to refine AI behaviour, such as setting parameters, have been shown to significantly enhance user satisfaction and the alignment of AI outputs with user expectations (Jain et al., 2023; Amershi et al., 2014; Holstein et al., 2019; Viswanathan et al., 2022), maintaining user engagement, and ensuring that AI systems serve the intended purposes (Ueno et al., 2022; Cai et al., 2019). Improvement of AI systems through user feedback represents the third crucial dimension of HCAI. Human-in-the-loop approaches, where users contribute to the iterative development and refinement of AI models, have gained traction as a means of optimising AI (Dudley and Kristensson, 2018; Amershi et al., 2014; Fails and Olsen, 2003). Involving end-users in the process, is not only about improving the AI but also about ensuring that the systems evolve in ways that are aligned with user needs and ethical considerations (Luchs, 2023). In summary, HCAI efforts (Capel and Brereton, 2023) can be categorised into three fundamental objectives: understanding, controlling, and improving AI systems. Borrowing from the cybernetic loop (Patten and Odum, 1981), akin to sensing, processing, and reacting; these categories form the basis for a continuous cycle of interaction between humans and algorithms. We plan to employ these three lenses as cues for our co-design of user-LLM interactions, ensuring that the development of AI systems remains aligned with the needs and expectations of everyday users (Wiener, [n. d.]; Bao et al., 2021).
The design of CUIs (Landay et al., 2019; Diederich et al., 2022; Clark et al., 2019) requires a deep understanding of HCI principles, ensuring that these systems can handle the nuances of human dialogue, such as turn-taking, repair mechanisms, and emotional engagement (Landay et al., 2019; Cruz-Sandoval and Favela, 2019; Kaye et al., 2018). As the field progresses, there is a growing emphasis on addressing these challenges to enhance the user experience and make conversational interactions more natural and effective (Lee et al., 2019; Hwang et al., 2019). The emergence of large language models (LLMs) has revolutionised user interactions across various domains, providing unprecedented capabilities in understanding and generating human language. In healthcare, LLMs have been integrated into administrative and clinical tasks, streamlining operations such as patient record management, appointment scheduling, and even aiding in medical diagnostics, thereby enhancing efficiency and reducing the burden on healthcare professionals (Gebreab et al., 2024; Dao et al., 2024; Brown et al., 2020). One of the most promising applications of LLMs is in the realm of personal wellbeing, where they are employed in the development of conversational agents designed to support mental health, offering personalised interactions that help users manage stress, anxiety, and other mental health conditions (Ferrara, 2022; Ke et al., 2024). Despite the promising advancements, issues such as hallucinations (Xu et al., 2024), task-specific limitations (Mizrahi et al., 2024), and user experience problems persist (Wang et al., 2024). Innovative human-centred solutions, are needed to build reliable, interactive and creative agents tailored to unique needs (Shankar et al., 2024). In the context of parenting + AI, this calls for a closer examination of how AI can offer useful and usable support while addressing the unique challenges in this domain.

### v2

@@ Abstract.
Parenting brings emotional and physical challenges, from balancing work, childcare, and finances to coping with exhaustion and limited personal time. Yet, one in three parents never seek support. AI systems potentially offer stigma-free, accessible, and affordable solutions. Yet, user adoption often fails due to issues with explainability and reliability. To see if these issues could be solved using a co-design approach, we developed and tested NurtureBot, a wellbeing support assistant for new parents. 32 parents co-designed the system through Asynchronous Remote Communities method, identifying the key challenge as achieving a “successful chat.” As part of co-design, parents role-played as NurtureBot, rewriting its dialogues to improve user understanding, control, and outcomes. The refined prototype, featuring an Interaction Layer, was evaluated by 32 initial and 46 new parents, showing improved user experience and usability, with final CUQ score of 91.3/100, demonstrating successful interaction patterns. Our process revealed useful interaction design lessons for effective AI parenting support.
Parental Wellbeing, Perinatal Support, LLMs, Human-Centred AI, Interaction Design
journalyear: 2025conference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3714088isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Empirical studies in interaction designccs: Computing methodologies Intelligent agentsccs: Human-centered computing HCI theory, concepts and modelsccs: Applied computing Health care information systems
Figure 1. Overview of the four-part study detailing the iterative development and testing of NurtureBot, a parental wellbeing assistant.
@@
A visual representation of the four-part iterative development of NurtureBot involving two parent cohorts. The chart shows the progression from the initial prototype to NurtureBot v1, followed by ARC co-design sessions, NurtureBot v2, and NurtureBot v3. Key findings include conversation starters, tone, parenting resources, and interaction improvements. NurtureBot v1 identified seven major problems, leading to an interaction layer for NurtureBot v2, which addressed four feature requests. NurtureBot v3 showed further improvements with five new feature requests and achieved the highest CUQ score of 91.3.
@@ 1. Introduction
Parenting, often hailed as the most important job in our world, exacts a significant toll on both mind and body. The perinatal period represents a critical developmental stage for the child, the parent(s), and the formation of attachment relationships (Li et al., 2017; Allan et al., 2013). This period, spanning from the onset of pregnancy to the first year postpartum, is marked by profound physical, emotional, and psychological changes, which significantly impact parental wellbeing. In particular, many expectant and new mothers encounter substantial challenges, in mental health, emotional adjustment, and caregiving (Widarsson et al., 2012), in addition to the transformation their body goes through. During this period, parents experience heightened stress, anxiety, and feelings of isolation, but only one in three parents seek help (UNICEF UK, 2022). Perinatal care literature identifies several key issues, including inadequate access to mental health resources, availability of healthcare professionals, cost, stigma surrounding parental mental health, and challenges in identifying and supporting at-risk mothers (Rallis et al., 2014; Davis et al., 2022; Howard and Khalifeh, 2020). These challenges have been further aggravated by the COVID-19 pandemic (Bridle et al., 2022; Kinser et al., 2022). As a result, we see a growing body of research exploring how digital technologies can enhance perinatal care (van den Heuvel et al., 2018; Ginja et al., 2018).
@@
Notable Human-Computer Interaction (HCI) research has laid the groundwork for exploring how digital interventions can increase wellbeing across aspects such as physical and mental health of a family (Pina et al., 2017), a child’s growth (Kientz et al., 2009), sleep tracking of parents and children (Pina et al., 2020), and managing family screen-time behaviours (Hiniker et al., 2016). Today, with the availability of Artificial Intelligence (AI) to everyday users, the use of general generative applications such as ChatGPT for parenting is on the rise (Quan et al., 2024). With responsible design, Large Language Models (LLMs) have the potential to facilitate a range of services, from offering self-guided mental health support (Sharma et al., 2024), real-time emotional support (Kang et al., 2024) to providing evidence-based information at scale (Lai et al., 2023). The current use of LLMs to support parental and child wellbeing raises both optimism, but also ethical caution (Ashraf, 2024; Stade et al., 2024), due to limitations and uncertainties regarding accuracy of information provided, empathy, and usefulness (Kaneda et al., 2023).
@@
Given these challenges, our research adopts an exploratory approach, probing into a small subset of these broader issues. We focus on understanding how interactions between LLMs and parents might be designed to be helpful and engaging. Our goal is to explore what constitutes a usable and useful interaction in these contexts and how LLMs can be integrated into wellbeing support in ways that are practically engaging and thus useful for parents. To this end, we conducted a four-part online study involving 78 parents, who trialled iterative improvements of NurtureBot, a technology probe, designed to support parental wellbeing. This study, while not aimed at delivering a finalised solution, served as an exploratory step toward understanding how parents interact with LLM-based agents and identifying necessary features to make these tools genuinely supportive. The following Research Questions (RQs) guide our investigation into the user interaction aspects of engaging with an LLM-based parental wellbeing assistant:
@@
This four-part user-centred design process, as illustrated in Figure 1, guided the iterative development of NurtureBot. To mitigate emotional risks, we engaged a ‘slightly-out-of-target’ group of parents (Islind et al., 2023), each with a child aged five or six. These parents simulated stages of their perinatal period and reflected on past challenges with NurtureBot. Using a mixed-methods approach, we gathered both qualitative and quantitative feedback by testing NurtureBot v1, v2, and v3, each over a five-day period. On the final day of each testing period, participants completed the Chatbot Usability Questionnaire (CUQ) (Holmes et al., 2019), assessing usability and engagement (see Figure 2).
@@
  • Part I: The Initial Prototype: NurtureBot v1 The development of NurtureBot v1 leveraged LLM technology, informed by a previous community wellbeing initiative run by Committee For Children (Committee for Children, [n. d.]) to support new parents. Zero-shot prompting was used to integrate supportive conversation starters and key features, including empathetic chatting, wellbeing exercises, and parenting information. A cohort of 32 parents engaged with NurtureBot v1, generating 144 conversations. Thematic analysis identified seven key issues faced by the user. NurtureBot v1 achieved a CUQ score of 85.4, surpassing the benchmark of 68 (Holmes et al., 2019), indicating good usability.
@@
  • Part II: ARC Co-Design Study The Asynchronous Remote Community (ARC) (MacLeod et al., 2016; Prabhakar et al., 2017) co-design sessions engaged the same cohort of parents with Miro board activities (Miro, 2011). Parents prioritised the key issue as having a “successful chat,” and reimagined and rewrote NurtureBot’s dialogues to improve user understanding, control, and outcomes; thematic analysis of revised dialogues yielded seven refined user needs.
@@
  • Part III: The Interaction Layer: NurtureBot v2 NurtureBot v2 introduced an interaction layer with four essential components (See Figure 6): (i) A core interaction principle that enabled users to ”Understand, Control, and Improve” exchanges, (ii) Interaction states that directed the chatbot’s behaviour across various tasks initiated by the user, (iii) Interaction levels based on few-shot prompting, incorporating co-designed dialogues from ARC sessions, and (iv) Interaction elements, tailored to address user needs identified during ARC. NurtureBot v2 underwent testing with the same initial cohort, yielding 104 conversations, revealing some remaining minor bugs, while achieving a CUQ score of 87.4, indicating further improvement.
@@
  • Part IV: Prototype Improvement: NurtureBot v3 and Validation To ensure an unbiased evaluation, we refined NurtureBot further and tested v3 with a new cohort of 46 parents. Feedback analysis identified two new problems to be considered for future iterations, and a CUQ score of 91.3, marking a robust improvement in usability.
@@
This study initiates our exploration into co-designing agents for parental wellbeing, setting foundations for future research to refine and expand AI tools. The diverse parenting experiences and emotional complexities highlight the need for continued development and testing.
@@ 2.1. Supporting Family Wellbeing with Digital Tools
Within the field of HCI, there has been a long-term research agenda focusing on the study and development of impactful digital tools that offer a variety of social, physical, and mental wellbeing support to parents and mothers specifically (D’Ignazio et al., 2016; Thomas et al., 2019; Britton et al., 2019). The design and evaluation of digital tools aimed at promoting maternal wellbeing during the transition to motherhood (Newhouse, 2019), bump2bump, emphasise the need for holistic approaches in HCI that consider the challenges faced by new mothers. The role of peer support in digital environments is highlighted by the finding that online networks can provide valuable emotional and practical assistance to parents, particularly in situations where traditional face-to-face support may not be available (Yamashita et al., 2022). While apps supporting first-time mothers, such as Baby Buddy in the UK (Deave et al., 2019), are generally well-received, effectiveness depends on alignment between usability and the mothers’ understanding of their needs, suggesting a need to involve parents in the design process to ensure that digital tools provide meaningful support.
@@
The integration of AI chatbots into parenting and family well-being has opened new avenues in this context (Yu et al., 2023; Entenberg et al., 2021; Wong et al., 2021). Recent studies have focused on parenting with conversational agents such as Alexa (Beneteau et al., 2020), Siri (Sezgin et al., 2022) and other voice-activated agents in the home (Sun et al., 2021). The integration of LLMs into parenting practices (Quan et al., 2024), presents new possibilities and challenges, with LLM-based agents such as ChatGPT being used by parents of young children (Quan et al., 2024; Kaneda et al., 2023), though their impact on parenting remains largely unexplored. The performance and acceptance of ChatGPT’s responses in the childcare field (Kaneda et al., 2023), is also being studied with caution. Despite technological advancements, several limitations persist such as misinformation (Kaneda et al., 2023), lack of diversity (Bajwa et al., 2024), and threat of data misuse (Neerincx et al., 2021). While gamification can be an effective motivational tool (Wernbacher et al., 2022), it does not appeal to all users equally. The need for humane interaction with a good user experience has been repeatedly reported in the literature for digital health tools built with or without AI (Bautista and Schueller, 2023). A review of 13 AI-based mental health apps with user ratings over 4 out of 5, revealed limited support for AI literacy and explainability (Alotaibi and Sas, 2023). A study of engagement with 93 health apps indicates that while install rates are high, sustained user engagement remains low, with the median 15- and 30-day retention a mere 3.9% and 3.3% respectively (Baumel et al., 2019). These gaps suggest a need for further research and development of applications which can deliver sound AI-powered digital tools for wellbeing in a sustained way.
@@ 2.2. Interacting with AI for Everyday Users
As AI permeates daily life, from recommendation systems to digital assistants, the ability of non-expert users to engage meaningfully with AI systems becomes critical (Eiband et al., 2019). While Explainable AI (XAI) tools enable expert users to grasp the rationale behind AI outputs, (Gunning et al., 2019; Miller, 2023a; Carvalho et al., 2019), most non-experts do not engage with XAI (Miller, 2023b), highlighting the need for innovative methods to foster AI understanding (Miller, 2023b). Control over AI systems by everyday users is another essential dimension of Human-Centred AI (HCAI), with research underscoring the importance of user agency in human-AI interactions and challenges with it (König, 2024). Interactive interfaces that allow users to refine AI behaviour, such as setting parameters, have been shown to significantly enhance user satisfaction and the alignment of AI outputs with user expectations (Jain et al., 2023; Amershi et al., 2014; Holstein et al., 2019; Viswanathan et al., 2022a), maintaining user engagement, and ensuring that AI systems serve the intended purposes (Ueno et al., 2022; Cai et al., 2019). Improvement of AI systems through user feedback represents the third crucial dimension of HCAI. Involving end-users in the process, is not only about improving the AI but also about ensuring that the systems evolve in ways that are aligned with user needs and ethical considerations (Luchs, 2023). In summary, HCAI efforts (Capel and Brereton, 2023) can be categorised into three fundamental objectives: understanding, controlling, and improving AI systems. Borrowing from the cybernetic loop (Patten and Odum, 1981), akin to sensing, processing, and reacting; these categories form the basis for a continuous cycle of interaction between humans and algorithms. We plan to employ these three lenses as cues for our co-design of user-LLM interactions, ensuring that the development of AI systems remains aligned with the needs and expectations of everyday users (Wiener, [n. d.]; Bao et al., 2021).
The design of Conversational User Interfaces (Landay et al., 2019; Diederich et al., 2022; Clark et al., 2019) requires a deep understanding of HCI principles, ensuring that these systems can handle the nuances of human dialogue, such as turn-taking, repair mechanisms, and emotional engagement (Landay et al., 2019; Cruz-Sandoval and Favela, 2019; Kaye et al., 2018). As the field progresses, there is a growing emphasis on addressing these challenges to enhance the user experience and make conversational interactions more natural and effective (Lee et al., 2019; Hwang et al., 2019).
The advent of LLMs has propelled CUIs into a wide range of domains. In healthcare, LLMs have been integrated into administrative and clinical tasks, purporting to streamline operations such as patient record management, appointment scheduling, and even aiding in medical diagnostics, in the name of enhancing efficiency and reducing the burden on healthcare professionals (Gebreab et al., 2024; Dao et al., 2024; Brown et al., 2020). Another potential application of LLMs is in personal wellbeing, with agents offering personalised interactions that aim to help users manage stress, anxiety, and other mental health conditions (Ferrara, 2022; Ke et al., 2024). A systematic review (Abd-Alrazaq et al., 2020) of chatbots in improving mental health outcomes, concluded that chatbots might effectively reduce symptoms of depression, anxiety, and stress, while enhancing well-being and quality of life. However, the authors emphasised the necessity for rigorous evaluation of these tools to ensure their safety and efficacy.
ChaCha (Seo et al., 2024), a chatbot leveraging LLMs was developed to encourage children to share emotions related to personal events, demonstrating the potential of LLMs in facilitating emotional expression among younger populations. Key findings underline the chatbot’s ability to facilitate empathetic, developmentally appropriate interactions, though challenges such as balancing engagement and over-reliance were noted. Similarly, MindfulDiary (Kim et al., 2024), an LLM-powered tool was designed to assist psychiatric patients in journaling, was tested in a four-week field study involving 28 patients and five psychiatrists, demonstrating that the app facilitated consistent journaling, aided self-reflection, and enhanced patient-clinician communication. This study highlights the potential of LLMs to bridge gaps in mental health care, while addressing challenges of engagement, safety, and integration into clinical workflows. Another recent work explored using LLMs to craft narratives about postpartum depression (Progga et al., 2024), generating 45 stories based on themes from online support communities. 85% of the generated narratives successfully adhered to prompts, demonstrating a potential for addressing perinatal mental health challenges through personalised storytelling. However, the stories were also reported to frequently fall short of capturing the emotional depth and relatability inherent in user-generated stories, leaning instead towards methodical and structured exposition.
Despite the promising advancements in LLMs within healthcare, issues such as hallucinations (Xu et al., 2024), task-specific limitations (Mizrahi et al., 2024), and user experience problems persist (Wang et al., 2024). The ethical considerations surrounding AI-driven wellness applications have been scrutinized, particularly regarding their potential health risks and regulatory challenges (Freitas and Cohen, 2024). While the studies described above suggest that generative AI chatbots can offer meaningful assistance to users in the wellbeing domain, concerns persist about their safety and effectiveness, necessitating further research to validate their use in unprompted, unguided real-world scenarios (Siddals et al., 2024). Adopting a more human-centred approach may help to build reliable, interactive and creative agents tailored to unique needs (Shankar et al., 2024). In the context of parenting + AI, this calls for a closer examination of how AI can offer useful and usable support while addressing the unique challenges in this domain.